You drive 79 miles in 2 hours in a perfectly straight line.

What is your average velocity in kilometer per mile(round to the nearest 0.1 km/hour)?

(You may need to look up how many km = 1 mile.)
Answer #1

Total displacement travel = 79 miles

1 mile = 1.609 km

Total displacement in km = 79*1.609 = 127.1 km

Total time taken = 2 hours

Avg velocity = total displacement / total time taken

= 127.1/2 = 63.5 km/hour